# secReferences in the Online Appendix

### v5

@@ 4.1 Stakeholders’ Reactions
This section explores how campaign timing affects critical stakeholders of a targeted firm, such as shareholders and consumers, while Section 4.2 examines its ramifications for the target’s operations. The relevant exhibits are in the Online Appendix.
@@ 4.1.1 Shareholders
Shareholders can be seen as primary proponent of firm’s social values (e.g., Naaraayanan et al., 2021; Baker et al., 2022). Shareholders may respond more to causes campaigned during AGMs, when media coverage is highest. We define as “related” a shareholder proposal on the same cause of the campaign in the previous year. Appendix C describes the matching procedure.
For each campaign in our sample, we look at the probability that the next AGM has a related shareholder proposal, depending on the timing of the campaign:
[…]
$+\alpha_{i,s(n),g,y}+u_{i,n,y}$
[…]
Figure B7 shows that only campaigns launched twelve months before the next AGM increase the likelihood of a shareholder proposal on the same issue. The effect is substantial—an increase of 12 percentage points, or 44% relative to the baseline. Since SEC 14a8(e)(2) rule prevents shareholders from submitting proposals within 120 days of the proxy statement’s release, the x-axis is capped at four months before the previous year’s AGM. This timing suggests that campaigns shortly after an AGM may spur proposals at the next one, possibly due to heightened media attention.
Importantly, we can rule out the possibility that NGOs partnering with shareholders drive this outcome. If that were the case, submitting a shareholder proposal followed by an early campaign would allow the partnering shareholders and NGOs to immediately gain broad support on the shareholder’s proposal, without having to wait a year to advocate for change at the next AGM.
@@ 4.1.2 Consumers
$\frac{\text{Sales}_{n,q}}{\text{Assets}_{n,q-1}}$ $=\sum^{\tau=5}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\beta_{\tau}\cdot\text{NGO Campaign}_{n}\cdot\mathbf{1}{\{\text{distance}=\tau\}}\right)+\alpha_{n}+\alpha_{s(n),q}+u_{n,q},$ (3)
where $q$ indicates a quarter and $\alpha_{n}$ and $\alpha_{s(n),q}$ are fixed effects at either the firm level or the industry-by-quarter-by-year level, respectively. The dependent value normalizes $n$’s sales by its asset to account for time variation in relative firm sizes, which is large in our ten-year sample. Standard errors are clustered at the firm level.
Appendix Figure B8 shows the estimated ${\hat{\beta}\tau}\tau$ coefficients separately for NGO campaigns (i) planned on an AGM date (Panel a) and (ii) not planned on an AGM date (Panel b). We find no consumer response to the latter, but sales drop by 0.7pp (about 5%) for the former—a magnitude consistent with Conway and Boxell (2024); Saint-Jean (2024). In Panel (c), the sales gap between AGM and non-AGM campaigns emerges one quarter before the event and continues to widen, possibly due to the greater media coverage of AGM dates.
@@ 4.2 Does AGM-Targeting Create Positive Change?
If public shaming were an effective strategy to influence firm behavior, targeting the AGM month should yield the largest improvements in E&S scores. However, Appendix Figure B9 shows that only campaigns launched 3 to 6 months before the AGM lead to subsequent increases in firms’ E&S scores—rising by 2–3% over the following year. By contrast, campaigns launched at other times, including during the AGM month, have no economically or statistically significant effect.
Although ESG scores are an imperfect measure of a firm’s social impact (e.g., Kotsantonis and Serafeim, 2019; Allcott et al., 2022), this finding suggests that NGOs’ preference for targeting AGM dates may be driven by factors beyond their direct influence on firms. Therefore,
@@
NGO campaigns on an AGM date impact stakeholders but do not lead to improvements in E&S scores. (i) Shareholders are more likely to make related proposals the following year. (ii) The target’s sales drop in the months after a campaign.
5 Timing Campaigns to Influence AGM Votes
@@
$\text{Success}_{i,n,m,y}=\sum^{0}_{\begin{subarray}{c}l=-6\end{subarray}}\beta_{l}\cdot\text{Campaign}_{i,n,m+l,y}+\alpha_{n,y}+\alpha_{i,y}+\alpha_{m}+u_{i,n,m,y},$ (4)
where “$\text{Success}_{i,n,m,y}$” is an indicator for the success of proposal $i$ at firm $n$ in month $m$ in year $t$, and “$\text{Campaign}_{i,n,m+l,t}$” is an indicator for the occurrence of an NGO campaign on the same cause as proposal $i$, $l$ months before the AGM. The reference period covers the months that are more than six months before the AGM of year $y$—i.e., $m\in[-11,-7]$. We account for firm-by-year, and cause-by-year fixed effects, as well as seasonality with month fixed effects, thereby exploiting only variation across campaigns and causes within firm-year pairs. The standard errors are clustered by firm.
@@
(a) Proposals accepted or withdrawn
@@
Note: E&S denote environmental and social proposals, respectively. The figure plots the estimated coefficients from (4). The dependent variables are an indicator variable that is one if a shareholder proposal is withdrawn or successful if voted (911 proposals) in Panel (a), and the percentage of votes in favor of a proposal (2,078 proposals) in Panel (b). Standard errors are clustered at the firm level.
Figure 3(a) plots the estimated $\{\hat{\beta}_{l}\}_{l}$. We find that NGO campaigns happening 6 months and 1 month before the AGM positively influence the outcome of related shareholder proposals. They increase, respectively, the probability of success by 2pp (6%) and 3pp (8%) more than a campaign launched more six months earlier.
To understand why campaigns 6 months before the AGM are also somewhat successful, we estimate (4) on the subset of voted proposals, with the percentage of votes in favor as the dependent variable. Note that every proposal to be voted on at the AGM is accompanied by a vote recommendation by the management (“For”, “Against”, or “Abstain”). Institutional investors often follow proxy advisor recommendations, particularly from ISS (Shu, 2024). We include management and ISS recommendations as covariates in (4).
Figure 3(b) shows that only NGO campaigns launched one or three months before the AGM increase support for related shareholder proposals, with an increase of nearly 2 percentage points. Given the generally low baseline support for these proposals (24%), this corresponds to an 8% increase in shareholder support when an NGO campaign addresses the same cause, on average.
In contrast to Figure 3(a), Figure 3(b) detects no significant increase in support for campaigns initiated as early as six (or five) months before the AGM. This suggests that the difference between the two figures arises from shareholder proposals being withdrawn after reaching an agreement with the firm. As a result of these early campaigns, the target firm may take actions to resolve the issue highlighted in both the campaign and the shareholder proposal, leading to an improvement in its E&S score, which supports the finding in Section 4.2 that only early campaigns result in better E&S scores.
Moreover, the fact that campaigns launched one to three months before an AGM help secure more votes aligns with the timing of proxy voting. Shareholders typically receive voting materials within 90 days of the AGM, a period when AGM visibility in the news may increase as well as the salience of the campaign for shareholders (ISS, 2022).
Note that the subset of shareholder proposals going to a vote is not random; an argument raised, for instance, by Flammer (2015). For this reason, we investigate all E&S proposals in Figure 3(a), which also include withdrawn proposals. Given the low support of proposals going to vote, managers may let proposals that are the least likely to succeed go to a vote. However, those votes can still be helpful in disciplining firms as large shareholder support to a cause may be hard to conceal for managers (Levit and Malenko, 2011).
@@
NGO campaigns are more effective at influencing a firm’s decisions when launched before the AGM.
@@
Sections 4 and 5 outline a tradeoff: late campaigns expose a firm’s wrongdoing to its stakeholders and boost the NGO’s visibility, while early ones are more effective at directly influencing a firm’s decisions. How do NGOs navigate this tradeoff?
In this section, we first explore whether NGOs, after gaining visibility by targeting AGMs in the past, switch to campaigning before the AGM. We then examine whether this “visibility capital” from previous campaigns increases the likelihood of success at the AGM for shareholder proposals related to the NGO campaign.
6.1 Accumulating Visibility
Targeting AGM-dates could enahnce an NGO’s visibility and its reputation for commitment to a cause. In this section, we explore whether NGOs, after accumulating some visibility, begin launching campaigns earlier—before the target firm’s AGM—to improve their chances of influencing the outcome as suggested in Section 5.
We investigate this mechanism by:
@@ 6.1 Accumulating Visibility
$\text{Distance to AGM}_{c(i,n,g),y}$ $=\sum_{\begin{subarray}{c}\tau>1\end{subarray}}\beta_{\tau}\cdot\mathbf{1}{\{\text{Campaign \#}_{c(\cdot),g,y}=\tau\}}$ (5)
$+\alpha_{g,y}+\alpha_{n,y}+\alpha_{i,y}+u_{c(g,n,i),y},$
which examines “$\text{Distance to AGM}_{c(i,n,g),y}$,” the distance (in months) of campaign $i$ to the AGM of targeted firm $n$, as a function of the number of past campaigns by NGO $g$, using an indicator, “$\text{Campaign \#}_{c,g,y}$,” that is one if the number of NGO campaigns launched by NGO $g$ until year $y$ is equal to $\tau$ and zero otherwise.
Since our campaign data start in 2010, we do not know how many campaigns an NGO launched before 2010. To control for this truncation as well as time varying confounders such as NGO size, we include NGO-by-year fixed effects. We also include firm-by-year and cause-by-year fixed effects to account for firm and cause-specific occurrences that, varying over time, may incentivize an NGO to launch a campaign. Hence, the identification of $\{\beta_{\tau}\}_{\tau}$ comes from multiple campaigns $c(\cdot)$ by the same NGO $g$ within year $y$ [^11] In Appendix Figure B10, we present similar results when controlling only for firm and NGO fixed effects, so that the estimation of $\hat{\beta}_{\tau}$ come from variation across both campaigns and years. Because of the scope of this analysis, we only consider NGOs with at least two campaigns, but the results still hold in the full sample. The standard errors are clustered at the NGO level.
To interpret the estimated $\{\hat{\beta}_{\tau}\}_{\tau}$ coefficients, consider that, on average, the first NGO campaign is exactly on the month of the AGM. Thus, Figure 4 indicates that as the NGO’s experience increases, the new campaigns are launched before the AGM date. [^12] The large negative coefficients estimated for NGOs with many campaigns—e.g., more than 9—are driven by outlier NGOs with several campaigns per year. We postpone the interpretation of this result to Section 7, which sums up and discusses the results.
Alternative mechanism—learning. This result is unlikely due to NGOs learning that campaigning on the AGM date does not influence voting decisions, as most shareholders have already voted by proxy. Appendix Figure B11 shows the same downward trend observed in Figure 4 when subsetting to NGOs whose first campaign took place within 2 weeks after the AGM, when the NGO surely knew that it could not affect votes.
Alternative mechanism—selection. NGO do not increasingly launch campaigns against firms that may be more easily influenced over time. Appendix Figures B12 and B13 exclude selection on observables. For instance, target sizes do not covary (market capitalization) with NGO experience. In addition, environmental (social) scores do not covary with an NGO’s experience in environmental (social) campaigns.
@@
NGOs initially target AGM dates to gain visibility. They schedule subsequent campaigns a few months before AGM dates to influence the outcome of shareholder proposals.
Figure 4: More experienced NGOs campaign earlier
Note: This figure plots the estimated coefficients from (5). The dependent variable is the months between an NGO campaign and the targeted firm’s AGM date. Only firms with at least two campaigns are considered (1,136 observations). Standard errors are clustered at the NGO level.
6.2 NGO Success and Reputation
We investigate whether NGO reputation, as proxied by past campaigns on AGM dates, makes NGOs more effective at producing positive corporate change. We assess this channel with the following regression:
$\text{Success}_{i,n,y}$ $=\beta_{1}\,\text{Campaign on Same Cause}_{c(i,n,g),y}+\beta_{2}\,\text{Experience}_{g,y}$ (6)
$+\beta_{3}\,\text{Campaign on Same Cause}_{c(i,n,g),y}\cdot\text{Experience}_{g,y}$

### v6

@@ 4.1 Stakeholders’ Reactions
We next ask how campaign timing matters for stakeholders and firms’ social and environmental outcomes. We focus on shareholders and consumers, with supporting exhibits in the Online Appendix.
@@ 4.1.1 Shareholders
Shareholders can be seen as primary proponent of firm’s social values (e.g., Naaraayanan et al., 2021; Baker et al., 2022b). Shareholders may respond more to causes campaigned during AGMs, when media coverage is highest. We define as “related” a shareholder proposal on the same cause of the campaign in the previous year. Appendix D describes the matching procedure.
For each campaign in our sample, we examine the probability that the next AGM includes a related shareholder proposal, depending on the campaign’s timing. To this end, we recalendarize time so that month 0 corresponds to the AGM month for each firm. As discussed in Section 3, AGM months are stable over time; thus, this recentering does not alter the distribution of months across years or risk double counting. [^4] All subsequent analyses use the same recalendarization. Months –5 and +5 are defined as 45-day intervals to ensure the recalendarized year covers the entire calendar. We estimate the following regression:
$\text{Proposal on Same Cause}_{i,n,y}$ $=\sumop\displaylimits^{\tau=-4}_{\begin{subarray}{c}\tau=-15\\ \tau\neq-4\end{subarray}}\left(\beta_{\tau}\cdot\mathbf{1}_{\{\text{Distance to AGM${}_{c(i,n,g,y,m)}$}=-\tau\}}\right)$ (2)
$+\alpha_{i,s(n),g,m,y}+\gamma_{n}+u_{i,n,y}$
For each campaign $c(i,n,g,y,m)$ launched by NGO $g$ against firm $n$ on cause $i$ in year $y$ and month $m$, we analyze the likelihood that the next AGM of firm $n$ includes a shareholder proposal on the same cause. Formally, $\text{Proposal on Same Cause}_{i,n,y}$ equals one if such a proposal is observed within 4 to 15 months after campaign $c(\cdot)$, and zero otherwise, with timing measured by the indicator $\mathbf{1}_{\{\text{Distance to AGM}_{c(i,n,g,y,m)}=-\tau\}}$. Since SEC 14a8(e)(2) rule prevents shareholders from including proposals in the proxy statement within 120 days of its release, we cap the analysis at four months before the previous year’s AGM and set the reference month to $\tau=-4$ because of SEC rules.
Although firm-by-time fixed effects are not possible in this setting as they would absorb the variation used to estimate $\beta_{\tau}$, this analysis purges all the variation from cause-by-industry-by-NGO-by-month-by-year through $\alpha_{i,s(n),g,m,y}$ and firm invariant characteristics through $\gamma_{n}$, which we view as the main confounding factor since issue salience varies systematically across firms, industries, causes, and NGOs over time. We cluster standard errors by firm.
Figure B3 shows that only campaigns launched twelve months before the next AGM increase the likelihood of a shareholder proposal on the same issue. The effect is substantial—an increase of 12 percentage points, or 44% relative to the baseline. This timing suggests that campaigns around an AGM may spur proposals at the next one, possibly due to heightened media attention.
Importantly, we can rule out the possibility that NGOs partnering with shareholders drive this outcome. If that were the case, submitting a shareholder proposal followed by an early campaign would allow the partners to immediately gain broad support on the proposal, without having to wait a year to advocate for change at the next AGM.
@@ 4.1.2 Consumers
$\frac{\text{Sales}_{n,q}}{\text{Assets}_{n,q-1}}$ $=\sumop\displaylimits^{\tau=5}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\beta_{\tau}\cdot\text{NGO Campaign}_{n}\cdot\mathbf{1}_{\{\text{distance}=\tau\}}\right)+\alpha_{n,y}+\alpha_{s(n),q}+u_{n,q},$ (3)
where $q$ indicates a quarter and $\alpha_{n,y}$ and $\alpha_{s(n),q}$ are fixed effects at either the firm-by-year level or the industry-by-quarter-by-year level, respectively. The dependent value normalizes $n$’s sales by its asset to account for time variation in relative firm sizes, which is large in our ten-year sample. Standard errors are clustered at the firm level.
Appendix Figure B4 shows the estimated $\hat{\beta}_{\tau}$} coefficients separately for NGO campaigns (i) planned on an AGM date (Panel a) and (ii) not planned on an AGM date (Panel b). We find no consumer response in either case, indicating that shareholders may be the primary audience for NGO campaigns. [^5] Unreported results show also no difference between experienced and unexperienced NGOs.
@@ 4.2 Does AGM-Targeting Create Positive Change?
If public shaming effectively influenced firm behavior, the AGM month should deliver the largest improvements in E&S scores, given that it triggers heightened negative news coverage of the target firm (Appendix Figure B2). However, Appendix Figure B5 shows that only campaigns launched 3 to 4 months before the AGM lead to subsequent increases in firms’ E&S scores—rising by around 3% over the following year. By contrast, campaigns launched at other times, including during the AGM month, have no economically or statistically significant effect.
Although ESG scores are an imperfect measure of a firm’s social impact (e.g., Kotsantonis and Serafeim, 2019; Allcott et al., 2022), this finding suggests that NGOs’ preference for targeting AGM dates may be driven by factors beyond their direct influence on firms. Therefore, the results in this section suggest that:
@@
NGO campaigns on AGM dates spur shareholders to place related proposals at future AGMs of the targeted firm, but they produce no contemporaneous improvements in the firm’s E&S scores.
5 Timing Campaigns to Influence AGM Outcomes
@@
$\text{Success}_{i,n,m,y}=\sumop\displaylimits^{0}_{\begin{subarray}{c}\tau=-6\end{subarray}}\beta_{\tau}\cdot\text{Campaign}_{i,n,m+\tau,y}+\alpha_{n,y}+\alpha_{i,y}+\alpha_{m}+u_{i,n,m,y},$ (4)
where $\text{Success}_{i,n,m,y}$” is an indicator for the success of proposal $i$ at firm $n$ in month $m$ in year $t$, and $\text{Campaign}_{i,n,m+\tau,t}$” is an indicator for the occurrence of an NGO campaign on the same cause as proposal $i$, $\tau$ months before the AGM. The reference period covers the months that are more than six months before the AGM of year yy—i.e., $m\in[-11,-7]$, as campaigns launched during that period are more likely to be in reaction to the previous year’s AGM results. We account for firm-by-year, and cause-by-year fixed effects, as well as seasonality with month fixed effects, thereby exploiting only variation across campaigns and causes within firm-year pairs. The standard errors are clustered by firm.
@@
(a) Proposal accepted or withdrawn
@@
Note: The figure plots the estimated coefficients from (4). The dependent variables are an indicator variable that is one if a shareholder proposal is withdrawn or successful if voted (911 proposals) in Panel (a), and the percentage of votes in favor of a proposal (2,078 proposals) in Panel (b). Standard errors are clustered at the firm level.
Panel (a) Figure 3 plots the estimated ${\beta}_{\tau}$. We find that NGO campaigns happening 6 months and 1 month before the AGM positively influence the outcome of related shareholder proposals. They increase, respectively, the probability of success by 2 percentage points and 3 percentage points more than a campaign launched more than six months earlier, corresponding to a relative increase between 6 and 8%.
To understand why campaigns launched six months before the AGM are also somewhat successful, we re-estimate (4) on the subset of proposals that went to a vote, using the percentage of votes in favor as the dependent variable. [^6] Since this analysis focuses on voted proposals, we also control for the voting recommendations of management and proxy advisors such as ISS (“For,” “Against,” or “Abstain”). Panel (a) of Figure 3 shows that only NGO campaigns launched one or three months before the AGM increase support for related shareholder proposals, with an increase of nearly 2 percentage points. Given the generally low baseline support for these proposals (24%), this corresponds to an 8% increase in shareholder support when an NGO campaign addresses the same cause, on average.
In contrast to Panel (a), Panel (b) detects no significant increase in support for campaigns initiated as early as six (or five) months before the AGM. This suggests that the difference between the two figures arises from shareholder proposals being withdrawn after reaching an agreement with the firm. As a result of these early campaigns, the target firm may take actions to resolve the issue highlighted in both the campaign and the shareholder proposal, leading to an improvement in its E&S score, which supports the finding in Section 4.2 that only early campaigns result in better E&S scores.
Note that the subset of shareholder proposals going to a vote is not random; an argument raised, for instance, by Flammer (2015). Given the generally low support of shareholder proposals going to vote, managers may let proposals that are the least likely to succeed go to a vote. However, those votes can still be helpful in disciplining firms as large shareholder support to a cause may be hard to conceal for managers (Levit and Malenko, 2011). This reinforces the importance of analyzing all E&S proposals, including those withdrawn as in Panel (a).
The evidence in this section points to:
@@
Pre-AGM campaigns raise shareholder support and proposal success, while AGM-day campaigns have little effect.
@@
Sections 4 and 5 outlined a tradeoff: campaigns launched close to an AGM maximize exposure, while those launched earlier maximize influence. In this section, we examine how NGOs navigate this tradeoff as their visibility increases. We first study whether greater experience or reputation leads NGOs to shift campaign timing from the AGM date to earlier months. We then ask whether such reputation translates into greater effectiveness at influencing shareholder proposals.
6.1 NGO Experience and Campaign Timing
We first examine whether NGOs, as they gain experience, shift from targeting AGMs to launching earlier campaigns. Specifically, we estimate
@@ 6.1 NGO Experience and Campaign Timing
$\text{Distance to AGM}_{c(i,n,g,y,m)}$ $=\sumop\displaylimits_{\tau=0.1}^{0.9}\beta_{\tau}\cdot\mathbf{1}_{\left\{\text{Campaign \#}_{g,y-1}=\tau\right\}}$ (5)
$\quad+\alpha_{t(\cdot),y}+\alpha_{g,y}+\alpha_{n,y}+u_{c(i,n,g,y,m)}.$
where the dependent variable measures the number of months between campaign $c(\cdot)$ and the targeted firm’s AGM, and the key explanatory variable is the sequence number of campaigns by NGO $g$ by year $y$. $\alpha_{t(\cdot),y}$, $\alpha_{g,y}$, and $\alpha_{n,y}$ are topic-by-year, NGO-by-year, and firm-by-year fixed effects, respectively. Identification thus comes from NGOs launching multiple campaigns in the same year. [^7] Appendix Figure B6 shows similar results when controlling only for firm and NGO fixed effects, so that the estimation of $\hat{\beta}_{\tau}$ come from variation across both campaigns and years.
Panel (a) of Figure 4 shows that first campaigns tend to occur right at the AGM, while subsequent ones are launched progressively earlier. This pattern is consistent with NGOs initially using AGMs to gain visibility and later exploiting their visibility to influence voting outcomes.
These patterns are not driven by alternative explanations. First, they are unlikely to reflect NGOs learning that campaigning on the AGM date does not affect voting, since most shareholders have already voted by proxy. Indeed, the same downward trend in campaign timing holds when focusing on NGOs whose first campaign occurred after the AGM (Appendix Figure B7). Second, they are not driven by selection of more “influenceable” targets: neither target size (market capitalization) nor ESG scores covary with NGO experience (Appendix Figures B8 and B9).
6.2 Reputation and Campaign Timing
We next use Google Trends scores to proxy for an NGO’s reputation. We compute the average search score for NGO $g$ in year $y$ relative to Greenpeace in year y1y-1, discretized into deciles and regress:
$\text{Distance to AGM}_{c(i,n,g,y,m)}$ $=\sumop\displaylimits_{\tau=0.1}^{0.9}\left(\beta_{\tau}\cdot\mathbf{1}_{\left\{\text{Relative Google Score}_{g,y-1}=\tau\!\right\}}\right)$ (6)
$\quad+\alpha_{t(\cdot)}+\alpha_{y}+u_{c(i,n,g,y,m)},$
which is similar to model (6) with $\text{Relative Google Score}_{g,y-1}$ as a measure of NGO visibility. $\alpha_{t(\cdot)}$ and $\alpha_{y}$ are topic and year fixed effects. Panel (b) of Figure 4 shows that only the most visible NGOs—those in the top deciles of relative searches—campaign several months ahead of AGMs, while less visible NGOs continue to concentrate on AGM dates.
Taken together, these results indicate a systematic shift: NGOs start by targeting AGMs to maximize exposure, and as their reputation or experience grows, they move campaigns forward in time to increase their chances of influencing firm decisions.
@@ 6.2 Reputation and Campaign Timing
NGOs initially target AGM dates to gain visibility. As their experience or reputation increases, they advance campaigns by several months to influence the outcome of shareholder proposals.
Figure 4: NGO reputation and campaign timing
(a) More experienced NGOs campaign earlier
(b) More visible NGOs campaign earlier
Note: This figure plots the estimated coefficients from (5). The dependent variable is the months between an NGO campaign and the targeted firm’s AGM date. Standard errors are clustered at the NGO level.
6.3 NGO Reputation and AGM Influence
We now ask whether NGO experience and reputation, as measured in the previous sections, translates into greater influence over shareholder proposals. To test this, we estimate
$\text{Success}_{i,n,y}$ $=\beta_{1}\,\text{Campaign on Same Cause}_{c(i,n,g,y,m)}+\beta_{2}\,\text{Experience}_{g,y-1}$ (7)
$+\beta_{3}\,\text{Campaign on Same Cause}_{c(i,n,g,y,m)}\cdot\text{Experience}_{g,y-1}$